If a baseball (m=.450 kg) is hit with a force of 1700 N, what is the baseball's acceleration in m/s2?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52858)

.

To find acceleration/deceleration, divide the force by the mass, using consistent units.



Answer by math_tutor2020(3817)


Force = Mass*Acceleration
F = m*a
a = F/m
a = (1700 newtons)/(0.450 kg)
a = (1700/0.450) m/s^2
a = 3778 m/s^2
This result is approximate.